# Peri-LN: Revisiting Layer Normalization in the Transformer Architecture

### v2

@@ 3.4 Stability Analysis in Normalization Strategies
4 Experiments
In this section, we provide a comprehensive empirical comparison of Post-, Pre-, and Peri-Layer Normalization (LN) across large-scale Transformer pre-training and subsequent evaluations on the language domain.
@@ 4.1 Experimental Setting
Excluding the embedding parameters, model sizes were set to $400$M, $1.5$B, and $3.2$B parameters, respectively. Each model was trained on $30$ billion tokens. For a fair comparison, we pre-trained each model with five different training seeds. We conducted a systematic exploration of learning rates, ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$. The global batch size is set to $256$. We use Adam optimizer with cosine learning rate scheduler. The sequence length was set to $8192$, and the weight decay coefficient was fixed at $0.033$. We employed Megatron-LM [^4] https://github.com/NVIDIA/Megatron-LM to pre-train the Transformers under each layer normalization strategy. We used the DCLM-baseline dataset (Li et al., 2024a) for our experiments, along with the “cl100100100100k_base” version of the TikToken tokenizer [^5] https://github.com/openai/tiktoken. Unless otherwise noted, most training and model settings followed those of the DCLM experiments. For layer normalization, we primarily employed RMSNorm. Further details are provided in Appendix B.
@@ 4.2 Pre-Training Large Language Models
Figure 3(a) illustrates the pre-training loss across learning rates for models ranging in size from $400$M to $3.2$B parameters. Notably, the Peri-LN architecture consistently achieves superior loss curves over this entire model size. Since Pre-LN shows best performance at learning rate $2\times 10^{-3}$ across all model size, we set this to the default learning rate for Pre-LN and Peri-LN. Unlike Pre-LN, Post-LN’s appropriate learning rate lies in a lower range, so we have provided a separate summary in Appendix C.1. In Figures 3(b) and 3(c), we compare the pre-training loss and the gradient norm curve at each LN strategy’s best-performing learning rate of $400$M size models. The same trend was observed across different model sizes. In particular, when sweeping over training seeds, Pre-LN and Post-LN exhibited many spikes in the gradient norm curve, whereas Peri-LN showed relatively few spikes, supporting Proposition 3.1. This reduction is consistent across both large and small learning rates.
We provide additional results in Appendix F, including experiments using LayerNorm (instead of RMSNorm), analyses of varying sequence lengths, the effect of different training-token counts, and an ablation study on embedding LN. We also investigate pre-training using the OLMo2222-style Peri-LN architecture in Appendix G. These findings are likewise consistent with those previously reported.
@@
Figure 5: This figure shows the forward growth patterns of hidden states for different architectures, highlighting the structural impact of normalization placement. Each model contains $1.5$ billion parameters (excluding the embedding size). We confirmed that the observed trend remains consistent across all model sizes.
(a) Grad-norm at init
@@
(c) Grad-variance at init
@@
Figure 6: Comparison of the backward gradient norm and variance for Post-LN, Pre-LN, and Peri-LN Transformers at initialization and at the final stage. Model size is $1.5$B. We confirmed that the observed trend remained consistent across all model sizes. init denotes the initialization of the model parameters.
4.3 Benchmark Evaluations & Instruction Tuning
To evaluate how well the pre-trained model’s training loss aligns with its benchmark performance, we conducted five separate benchmark evaluations. Furthermore, to examine how supervised fine-tuning (SFT) improves scores and instruction-following capabilities under different layer normalization strategies in Transformer architectures, we performed additional training using the LIMA dataset (Ouyang et al., 2022; Zhou et al., 2023). Diverged checkpoints were excluded when calculating the evaluation score (mostly occurs in Pre-LN). Additional training hyperparameters for SFT are given in Appendix B.2. As shown in Table 2, Peri-LN consistently demonstrates superior performance across all model sizes. Additionally, we note that, beyond the improved scores, the standard deviation of the benchmark results across different training seeds is reduced by more than half with Peri-LN. From this, we observe that Peri-LN helps maintain consistency not only in gradient stability and final loss but also in benchmark performance. For the evaluation loss, we used $10$K random samples from the C4444 dataset (Raffel et al., 2020). Detailed configurations and individual scores are provided in Appendix H.
4.4 Systematic Analysis
Despite emerging empirical evidence that Peri-LN can stabilize training and bolster performance, many open questions remain: Which design factors are crucial for reaping its benefits at scale (§4.4.1)? How does it regulate hidden-state variance and gradient flow more effectively (§4.4.2, §4.4.3)? And why might certain variants outperform standard LN placements in large Transformer architectures (§4.4.4)? Addressing these points, we present a systematic analysis of Peri-LN’s mechanics at following sub-sections.
4.4.1 Learnable Parameter $\gamma$ of RMSNorm
To investigate the impact of module output normalization on training stability, as proposed in the Proposition 3.1, we fixed the learnable parameter $\gamma$ of RMSNorm to $1$, isolating the effect of normalization. As shown in Figure 4, we observe that simply normalizing the output of each module reduces gradient spikes and provides improvements in the loss comparing to the pre-normalization only transformers. Nonetheless, we also confirmed that allowing $\gamma$ to be learnable yields slightly better performance. We confirm that this trend remains consistent in both the $400$M and $1.5$B models. In this experiment, we omitted Peri-LN’s embedding layer normalization in order to isolate and evaluate the precise role and benefits of output layer normalization.
(a) At initialization
@@ 4.4.1 Learnable Parameter $\gamma$ of RMSNorm
(c) After $30$B tokens training
(d) Frozen $\gamma$ in Output-LN
Figure 7: Angular distance of hidden state is presented in Figure  7(a)7(c), and 7(d). In Figure 7(b), we monitor $\gamma$ of every Output-LN in Peri-LN during training. We use $30$B tokens trained $400$M size model in this experiments.
4.4.2 Growth of Hidden State
To examine in greater depth how Peri-LN affects the forward propagation, we analyzed the absolute magnitude and variance of the hidden states using $1,000$ samples from the Wikitext dataset (Merity et al., 2016). Figure 5 illustrates how normalization strategies influence the forward-path hidden states over the course of training.
Two particular aspects of hidden-state behavior were noted: First, with respect to model depth, Post-LN maintains stable hidden-state magnitude and variance owing to the presence of a normalization layer in the main path (“Highway”). By contrast, because Pre-LN does not normalize the outputs of each Attention and MLP module, the magnitude and variance of the hidden states exhibit exponential growth following the addition operation in the residual path. For the Peri-LN architecture, which leverages Output-LN, we observed that hidden states remain comparatively well-managed. Next, regarding changes over training iterations, Post-LN continues to normalize each block’s output, which prevents any substantial increase or decrease in the trend as training progresses. Meanwhile, Pre-LN exhibits a relatively linear variance distribution at initialization, but escalates exponentially to extremely large values over successive iterations. In contrast, Peri-LN fluctuates more moderately, owing to the role of Output-LN in controlling hidden-state magnitude and variance throughout training.
4.4.3 Layer-wise Gradient Norm & Variance
@@
4.4.4 Hidden State Representation
We utilize angular distance (Li et al., 2024b) as a metric to assess redundancy between hidden states at initialization and after training. This approach allows us to quantify how similar or distinct the representations were across layers. As shown in Figure 7(a), Post-LN exhibits smaller angular distances due to the LN being located on the main path, whereas Pre-LN and Peri-LN begin with very similar states. As shown in Figure 7(c), at the end of training, Pre-LN tends to produce more redundant hidden state representations compared to the others. This phenomenon may stem from Pre-LN’s repeated residual additions, which amplify certain representations over others.
To investigate this further, we focus on module output normalization, which is the main distinguishing factor between Pre-LN and Peri-LN. As shown in Figure 7(b), the learnable scale starts around $1$ in the early stages of training and gradually changes with increasing depth. Because Peri-LN preserves the identity path, it appears to adjust the scale of the module output accordingly. This suggests that the exponential growth of the main path’s magnitude in Pre-LN diminishes the relative contribution of individual modules, resulting in more redundant hidden representations. To support this, Figure 7(d) shows that fixing the learnable scale of Peri-LN’s module output LN at $1$ causes the main path contribution to decrease in deeper layers. This finding confirms the role of module output normalization in controlling hidden state redundancy.
@@ 5 Conclusion
We explore the placement of layer normalization within the Transformer architecture to better understand its role during training. By systematically comparing Post-LN, Pre-LN, and Peri-LN, we highlight their distinct impacts on stability, gradient propagation, and optimization dynamics. Our findings suggest that placing LN on module outputs in addition to the Pre-LN can help manage large activations while preserving beneficial gradient flow, thereby offering a promising balance for stable optimization. By unifying these approaches under the term Peri-LN, we seek to consolidate existing variants and encourage deeper investigation into this underexplored alternative.
@@ Acknowledgements
We thank our colleague Jeongin Bae for inspiring the underlying motivation for this research. We are also grateful to Jung Hyun Lee, Seonghyeon Kim, and Seunghyun Seo for their valuable assistance during the early stages of our research discussions. Finally, we extend our gratitude to Gichang Lee, Lead of the Backbone Mission at NAVER Cloud, for his unwavering support in connecting us with key resources throughout this research.
@@
Another positive externality of improving the cost-effectiveness of large-scale training is the reduction in the environmental impact of AI research. Training state-of-the-art LLMs requires vast computational resources, leading to significant energy consumption and carbon emissions. By enhancing the efficiency of large-scale training, our approach helps mitigate these environmental costs, making AI research more sustainable. Lowering the energy footprint of training not only aligns with global efforts toward greener computing but also enables a wider range of institutions, including smaller research labs and academic groups, to participate in cutting-edge AI development without excessive resource constraints. This shift promotes a more inclusive and responsible AI ecosystem, ensuring that advancements in foundation models are both accessible and environmentally conscious.
@@ Appendix A Related Work
Training stability in Transformer architectures is closely tied to the choice and placement of layer normalization (LN). Xiong et al. (2020) reported that Transformers employing a Post-LN structure often suffer from gradient instabilities at initialization, requiring a careful learning-rate warm-up phase to mitigate these issues. In contrast, Pre-LN helps maintain more stable gradients during the early stages of training. However, Kedia et al. (2024) showed that while Post-LN can lead to vanishing or exploding gradients in deep Transformers, Pre-LN may induce hyperbolic gradient growth. These findings illustrate the nuanced trade-offs of normalization placement and draw parallels to earlier CNN studies, where careful management of activation variance proved essential for stable deep learning.
@@
Appendix B Detailed Experimental Setting
B.1 Configurations on Pre-Training
Global Batch Size Weight Decay Iterations Optimizer LR Schedule Warmup
$256$ $0.033$ $14400$ Adam Cosine $10$%
Table 4: Embedding configurations.
$8192$ Rope $True$
Table 5: Model configurations.
Size $n_{layers}$ $n_{heads}$ $d_{model}$ $d_{head}$
$400$M $24$ $8$ $1024$ $128$
$1.5$B $24$ $16$ $2048$ $128$
$3.2$B $32$ $16$ $2560$ $160$
B.2 Configurations on Supervised Fine-Tuning
Appendix C Additional Results on Pre-Training Study
C.1 Post-Layer Normalization Architecture & Learning Rate Exploration
Figure 8: Learning rate explorations for Post-LN architectures.
C.2 Best Performing Checkpoints Comparisons of Other Model Sizes
(a) Training loss for $3.2$B
(b) Gradient-norm for $3.2$B
(c) Training loss for $1.5$B
(d) Gradient-norm for $1.5$B
C.3 Gradient Spikes during Pre-Training
Although this observation is empirical, it remains sufficiently noteworthy to discuss. During extensive pre-training ablation experiments, we frequently observed gradient-norm spikes. Specifically, when examining how these spikes vary with different learning rates, we found that Pre-LN exhibited irregular spikes regardless of the learning rate setting, with a particularly high occurrence in the early stages of training. In contrast, such spikes were rarely observed under Peri-LN.
Appendix D Proof of Theoretical Insight
To support the claim that Peri-LN enhances the stability of training in such cases, we analyze the gradient norm in the final layer. For simplicity, we use RMSNorm and ReLU here.
Proposition D.1.

### v3

@@ 4.1 Experimental Setting
Excluding the embedding parameters, the model size is set to the parameters $400$M, $1.5$B and $3.2$B, respectively. Each model is trained on $30$ billion tokens. To ensure reliable validation, we pre-train each model with five different training seeds in all experiments. We perform a exploration of the learning rates, ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$ to identify the U-shaped pattern for each LN strategy. The sequence length is set to $8192$, and the weight decay coefficient is fixed at $0.033$. We employ Megatron-LM [^4] https://github.com/NVIDIA/Megatron-LM to pre-train the Transformers under each LN strategy. We use the DCLM-baseline dataset (Li et al., 2024a), along with the “cl100100100100k_base” version of the TikToken tokenizer [^5] https://github.com/openai/tiktoken. Unless otherwise noted, most training and model configurations follow those of the DCLM experiments(Li et al., 2024a). For normalization layer, we primarily employ RMSNorm. Further details are in Appendix D.
@@ 4.2 Pre-Training Large Language Models
Figure 3(a) illustrates the pre-training loss across learning rates for models ranging in size from $400$M to $3.2$B parameters. Notably, the Peri-LN architecture consistently achieves superior loss curves over this entire model size. Since Pre-LN shows best performance at learning rate $2\times 10^{-3}$ across all model size, we set this to the default learning rate for Pre-LN and Peri-LN. Unlike Pre-LN, Post-LN’s appropriate learning rate lies in a lower range, so we provide a separate summary in Appendix E.1. In Figures 3(b) and 3(c), we compare the pre-training loss and the gradient norm curve at each LN strategy’s best-performing learning rate of $400$M size models. The same trend is observed across different model sizes (§E). In particular, when we sweep over training seeds and learning rates, Pre-LN frequently exhibits spikes in the gradient-norm curve, whereas Peri-LN shows comparatively few, thereby supporting Proposition 3.1.
(a) Gradient-norm at seed $5$
(b) Gradient-norm at seed $4$
Figure 5: Final-layer gradient norms for seeds $4$ and $5$.
4.3 Early Stage Instability in Pre-Training
Early in pre-training, Pre-LN models consistently show gradient spikes, loss surges, and occasional divergence across seeds and scales (Fig. 4). These issues are far less pronounced in Peri-LN. We posit that the instability of Pre-LN arises from three factors: (1) the hidden state variance exhibits a sudden surge from initialization through the early stages of optimization, deviating from the linear trend predicted by Eq. 4 (see §3.3); (2) the exponential growth of hidden state variance across both depth and training steps; and (3) the instability caused by the massive activations (Proposition 3.1). Among these, we highlight the variance growth along the main path as the principal driver of the observed divergence. To corroborate this, Section 6 presents targeted experiments that manipulate weight decay and weight initialization schemes, demonstrating how curbing extreme variance mitigates the instability of each LN strategy. The curves in 4(a), 4(b), and 4(c) are from a $400$M model, whereas 4(d) corresponds to a $1.5$B model.
4.4 Gradient Norm of the Final-layer
Motivated by Proposition 3.1, we track the final-layer gradient-norm in two representative runs selected from five training seeds. Figure 5(a), now including the newly added Peri-LN results, confirms the hierarchy reported by Xiong et al. (2020): whenever training remains stable, the gradient norms satisfy Post-LN $>$ Pre-LN $>$ Peri-LN. However, Figure 5(b) shows a run in which the Pre-LN model diverges even though every hyperparameter matches the stable run except for the random seed. In Section 5.1, we examine this failure in greater depth and relate it to Proposition 3.1. The curves in Figure 5 are obtained from $400$M models.
4.5 Benchmark Evaluations & Instruction Tuning
To evaluate how well the pre-training loss aligns with its benchmark performance, we conduct five separate benchmark evaluations. Furthermore, to investigate instruction-following capabilities under different layer normalization strategies, we conduct additional training using the LIMA dataset (Ouyang et al., 2022; Zhou et al., 2023). Diverged checkpoints are excluded from the evaluation score computation (mostly occurs in Pre-LN). Additional training hyperparameters for SFT are given in Appendix D.2. As shown in Table 1, Peri-LN consistently demonstrates superior performance across all model sizes. Additionally, we note that, beyond the improved scores, the standard deviation of the benchmark results across different training seeds is reduced by more than half with Peri-LN. From this, we observe that Peri-LN helps maintain consistency not only in gradient stability and final loss but also in benchmark performance. For the evaluation loss, we used $10$K random samples from the C4444 dataset (Raffel et al., 2020). Detailed settings and individual benchmark scores are provided in Appendix J.
@@ 4.5 Benchmark Evaluations & Instruction Tuning
Figure 6: Forward hidden state growth patterns for each LN strategy in a $1.5$B-parameter Transformer.
(a) Grad-norm at init.
@@
(c) Grad-variance at init.
@@
Figure 7: Backward gradient norm and variance of $1.5$B Post-, Pre-, and Peri-LN Transformers at initialization (init.) and final training.
5 Analysis
Despite emerging evidence that Peri-LN outperforms Post- and Pre-LN, key uncertainties remain: How do different LN placements shape hidden-state statistics and gradient flow (§5.1, §5.2)? What role does the Output-LN scale parameter $\gamma$ play (§5.3)? And why does Peri-LN produce more distinctive representations than its counterparts (§5.4)? The subsections below tackle these questions in turn.
5.1 Growth of Hidden State
To examine in greater depth how Peri-LN affects forward propagation, we analyze the absolute magnitude and variance of the hidden states using $1,000$ samples from the Wikitext dataset (Merity et al., 2016). Figure 6 shows how different normalization strategies influence forward-path hidden states over the course of training and across model depth. We observe the same pattern across all models trained with five different random seeds (§F).
Across layers, Post-LN maintains stable hidden state magnitudes and variances because the main path includes a normalization layer. In contrast, Pre-LN omits normalization after each attention and MLP sub-layer, so the magnitude and variance of the hidden states grow exponentially after the residual addition. For Peri-LN, which adds an Output-LN, these statistics remain comparatively well controlled. Across training iterations, Post-LN’s block-level normalization continues to suppress large shifts, preventing substantial drift in magnitude or variance. Pre-LN starts with an approximately linear variance profile at initialization but escalates exponentially to extremely large values as optimization proceeds. Peri-LN again exhibits only moderate fluctuations, owing to Output-LN’s consistent regulation of hidden-state statistics. Further discussion appears in Section 6.
(a) Training loss
(b) Loss in the final $5$B token interval
(c) Gradient-norm
(a) After $30$B tokens training
@@ 5.1 Growth of Hidden State
Figure 9: Angular distance between hidden states after training. Fig. 9(b) monitor $\gamma$ of every Output-LN in Peri-LN during training.
5.2 Layer-wise Gradient Norm & Variance
@@
5.3 Learnable Parameter $\gamma$ of RMSNorm
To investigate the impact of module output normalization on training stability, as proposed in the Proposition 3.1, we fix the learnable parameter $\gamma$ of RMSNorm to $1$, isolating the effect of normalization. As illustrated in Figure 8, adding output normalization to each sub-layer suppresses gradient spikes and lowers the loss relative to Transformers that employ only pre-normalization. Nonetheless, we also confirm that allowing $\gamma$ to be learnable yields slightly better performance. The trend persists consistently across model scales and random seeds. In this experiment, we omit Peri-LN’s embedding layer normalization in order to isolate and evaluate the precise role and benefits of output-LN.
5.4 Hidden State Representation
To assess hidden state redundancy after training, we employ angular distance (Li et al., 2024b), which quantifies how similar or distinct the layer representations are. As Figure 9(a) illustrates, Pre-LN produces markedly more redundant hidden states than the other variants by the end of training. We attribute this effect to the exponential growth of the main residual path in Pre-LN, which diminishes the relative contribution of individual sub-layers. In contrast, Peri-LN retains an identity path whose learnable scale begins near 1 and gradually adjusts with depth (Figure 9(b)), thereby moderating redundancy. These observations highlight the role of module-output normalization in controlling hidden state similarity. All statistics are computed on $256$ random samples from RedPajama-Data-1T (Computer, 2023). Appendix L includes additional figures and initialization comparisons.
6 Ablation Study
(a) Variance growth by weight decay
(b) Training loss by weight decay
(c) Variance growth by weight init.
Figure 10: Effects of weight decay and initialization (init.) on massive activations. 10(a) Strong weight decay relieves the variance explosion in Pre-LN. 10(b) Strong weight decay ($0.33$, which is 10×10$\times$10 × the baseline.) suppresses Pre-LN divergence. 10(c) Smaller-scale initialization slightly curbs Pre-LN variance, while Peri-LN remains bounded regardless. $d$ denotes the model’s hidden dimensionality.
To probe massive activations across conditions, we sweep weight decay coefficient and initialization variance for both Pre- and Peri-LN models, holding other settings fixed. Per-run results and detailed settings are in Appendix G.8.
6.1 Weight Decay
In Figure 10(a), stronger L2222 regularization markedly lowers the variance curve, confirming that heavier weight decay directly curbs forward-path explosions in Pre-LN. In contrast, the same increase in weight decay reduces Peri-LN’s variance growth only marginally. We take the stable run initialized with seed $3$ (Table 11) and sweep the weight decay coefficient. Table 7 in the Appendix further shows that, irrespective of the presence of massive activations, Peri-LN achieves better performance than Pre-LN.
To further probe stability under varying degrees of massive activation, we replicate the previously divergent run (seed $4$) and repeat the same weight decay sweep. As Figure 10(b) demonstrates, raising the weight decay coefficient from the baseline $0.033$ to $0.33$ (a tenfold rise) prevents divergence, providing empirical support for Proposition 3.1. Nevertheless, strong weight decay can stabilize Pre-LN, it still fails to close the performance gap relative to Peri-LN.
6.2 Weight Initialization
As the initialization variance increases, the severity of massive activations rises correspondingly for Pre-LN (Figure 10(c)); at the largest variance, the model diverges outright (Appendix G.8.3). Pre-LN therefore displays marked sensitivity to its initial conditions. In contrast, across the same range of ablations, Peri-LN’s loss curves and activation variances shift only marginally. This hyperparameter-insensitive robustness is corroborated by the low downstream standard deviations reported in Table 1.
6.3 Additional Results
For brevity, we defer an extensive set of supplementary experiments to the appendix. Appendix G reports the core robustness checks: replacing RMSNorm with LayerNorm, varying sequence lengths, reducing pre-training budgets, and ablating embedding-level normalization. OLMo2222-style Peri-LN pre-training runs appear in Appendix H. Stochastic gradient descent (SGD) baselines are summarized in Appendix I. Alternative LN placements, extending Figure 2, are provided in Appendix G.6. Across all settings, the results are consistent with the trends presented in the main Section 4 and 5, further substantiating our conclusions.
7 Implications
This section integrates our findings into practical guidance on variance-driven stability and precision constraints in large-scale Transformers.
7.1 Mitigating Variance-Driven Instability via Peri-LN
Pre-LN, the prevailing normalization strategy, is inherently prone to unchecked growth in activation variance (§5), which in turn induces numerical instability during training. Our extensive empirical analysis shows that Peri-LN— which normalizes the outputs of the Attention and MLP sub-layers—markedly curbs this variance and often prevents divergence (§4 & §6). Proposition 3.1 formalizes how excessive variance amplifies gradient norms, clarifying its causal role in destabilizing large-scale pre-training. In Pre-LN, instability is further exacerbated when the statistical conditions assumed at initialization depart markedly from those observed later in training (§3.3). By contrast, Peri-LN alleviates this discrepancy, thereby improving training stability, and delivering additional performance gains.
7.2 Precision Constraints Imposed by Pre-LN
Both Pre-LN and Peri-LN architectures leave the main hidden state path unnormalized, so once large values arise in earlier layers, they persist through to later layers. Consequently, Pre-LN’s additive residual path might generates activations near or beyond the FP16 limit. To gauge how often these values exceed FP16 yet remain within BF16, we track the top-$100$ absolute hidden state values for $3.2$B-parameter Pre-LN and Peri-LN models. In Figure 11, the blue band for Pre-LN surpasses the FP16 maximum bound as early as $0.5$B training tokens whereas Peri-LN (red band) consistently remains below this threshold. This pattern, echoing Sun et al. (2024), highlights that choosing FP16 or BF16 is not just a hardware preference but is closely linked to how hidden state magnitudes evolve within the model. Earlier work on OPT (Zhang et al., 2022), which was pre-trained using FP16 precision, suggests that the training instabilities they observed were likely exacerbated by numerical overflows and gradient pathologies (Proposition 3.1) arising when activations exceeded the representable range of FP16.
Figure 11: Evolution of extreme hidden state absolute magnitudes during training. Colored bands trace the range of the global top-$100$ absolute activations. Pre-LN (blue) quickly surpasses the FP16 representable maximum, while Peri-LN (red) stays below it.
@@ 8 Conclusion
We explore the placement of layer normalization within the Transformer architecture to better understand its role during training. By systematically comparing Post-LN, Pre-LN, and newly termed Peri-LN, we highlight their distinct impacts on stability, final performance, and optimization dynamics. Our findings suggest that placing LN on module outputs in addition to the Pre-LN can help manage large activations while preserving beneficial gradient flow, thereby offering a promising balance for stable optimization. By unifying these approaches under the term Peri-LN, we seek to consolidate existing variants and encourage deeper investigation into this underexplored alternative.
@@ Acknowledgements
We thank our colleague Jeongin Bae for inspiring the underlying motivation for this research. We are also grateful to Jung Hyun Lee, Seonghyeon Kim, and Seunghyun Seo for their valuable assistance during the early stages discussions. Finally, we extend our gratitude to Gichang Lee, Lead of the Backbone Mission at NAVER Cloud, for his unwavering support. This work was partly supported by Institute for Information & communications Technology Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST), No.RS-2021-II212068, Artificial Intelligence Innovation Hub, No. RS-2024-00509279, Global AI Frontier Lab)
@@
Improving the cost-effectiveness of large-scale training simultaneously lowers AI’s environmental footprint by reducing the vast energy consumption and carbon emissions inherent in state-of-the-art LLM development. This efficiency not only aligns with global sustainability goals but also enables smaller research labs and academic groups to pursue cutting-edge AI without prohibitive resource demands, fostering a more inclusive and responsible ecosystem.
@@ Appendix A Related Work
Training stability in Transformer architectures is closely tied to the choice and placement of layer normalization (LN). Xiong et al. (2020) reported that Transformers employing a Post-LN structure often suffer from gradient instabilities at initialization, requiring a careful learning-rate warm-up phase to mitigate these issues. In contrast, Pre-LN helps maintain more stable gradients during the early stages of training. However, Kedia et al. (2024) showed that while Post-LN can lead to vanishing or exploding gradients in deep Transformers, Pre-LN may induce hyperbolic gradient growth. These findings illustrate the nuanced trade-offs of normalization placement and draw parallels to earlier CNN studies, where careful management of activation variance proved essential for stable deep learning. Ding et al. (2021) introduced Sandwich-LN in the vision domain for the first time, yet they paid little attention to the structural characteristics and differences that distinguish one LN placement from another. In language domain, several major open-source language models (e.g., Olmo2222 (OLMo et al., 2024), Gemma2222 (Rivière et al., 2024), and Gemma3333 (Team et al., 2025)) already employ a Peri-LN-like structure (see Section 3). Nevertheless, previous technical reports have not explained why this design might be advantageous compared with the more widely studied Pre-LN and Post-LN. By investigating Peri-LN in detail, we aim to highlight the structural benefits that appear to underlie its empirical success in these implementations.
@@
Appendix B Proposition 3.1 of Post-LN
Proposition B.1.
Post-LN (vanishing gradient). Consider the following sequence of operations:
$a=\mathrm{MLP}(x),o=x+a,\tilde{o}=\mathrm{Norm}(o),$ (9)
then
where $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$. Since Post-LN normalizes the hidden state after each residual addition along the main path, the activation norm $|h|$ tends to remain relatively moderate. As a result, in Post-LN, the gradient scale is less sensitive to the magnitude of activations and more significantly influenced by model depth, as previously analyzed by Xiong et al. (2020) and Kedia et al. (2024), leading to vanishing gradients as depth increases.
Appendix C Proof of Theoretical Insight
To support the claim that Peri-LN enhances the stability of training in such cases, we analyze the gradient norm in the final layer. For simplicity, we use RMSNorm and ReLU here. Here, we assume that $\gamma$, the scaling parameter used in RMSNorm, is positive, and we empirically verified that it remains strictly positive across models of all sizes during training.
Proposition C.1.